# Can large language models generalize analogy solving like children can?

### v2

@@ Abstract
When we solve an analogy we transfer information from a known context to a new one through abstract rules and relational similarity. In people, the ability to solve analogies such as “body : feet :: table : ?” emerges in childhood, and appears to transfer easily to other domains, such as the visual domain “$(($ : )))) :: <<< : ?”. Recent research shows that large language models (LLMs) can solve various forms of analogies. However, can LLMs generalize analogy solving to new domains like people can? To investigate this, we had children, adults, and LLMs solve a series of letter-string analogies (e.g., a b : a c :: j k : ?) in the Latin alphabet, in a near transfer domain (Greek alphabet), and a far transfer domain (list of symbols). As expected, children and adults easily generalized their knowledge to unfamiliar domains, whereas LLMs did not. This key difference between human and AI performance is evidence that these LLMs still struggle with robust human-like analogical transfer.
Keywords analogy  $\cdot$ transfer  $\cdot$ generalization  $\cdot$ artificial intelligence  $\cdot$ reasoning
@@ 1 Introduction
You may be familiar with the analogy “consciousness is like an iceberg”. Here, people can intuitively infer the below-the-surface depth and complexity of consciousness by relating it to an iceberg, whose mass is mostly found under water, just as our subconscious dwells under our conscious minds. This intuitive ability emerges in childhood (Goddu et al.,, 2020; Gentner,, 1988; Stevenson and Hickendorff,, 2018). However, it is a subject of debate whether analogical reasoning has emerged in Large Language Models (LLMs) (Webb et al.,, 2023; Lewis and Mitchell,, 2024; Hodel and West,, 2023; Webb et al.,, 2024). More importantly, are LLMs able to solve analogies at this level of conceptual abstraction and generalize to novel domains (Mitchell,, 2021; Shiffrin and Mitchell,, 2023)? In this study, we investigate analogical transfer at two levels of abstraction (near and far), and compare LLM performance not only to adults, but also to children, who are still developing analogical reasoning abilities. We ask the question: Can LLMs can generalize analogy solving like children can?
@@
Analogical reasoning, the process of applying a known concept to understand something new through relational similarity, is fundamental to the way people think and learn (Holyoak,, 2012; Gentner and Hoyos,, 2017). This is because we humans can easily generalize – that is, transfer principles discovered in one domain to new domains that share varying degrees of similarity with the original (Doumas et al.,, 2022). This can be principles in near contexts that are similar in terms of concrete attributes (e.g., shape, “a pyramid is like an iceberg”) or in farther contexts that are only similar in terms of abstract relations (e.g., abstraction of depth, “consciousness is like an iceberg”) (Barnett and Ceci,, 2002). Near analogies tend to be easier for both adults and children to solve than far analogies (Stevenson et al.,, 2023; Jones et al.,, 2022; Thibaut and French,, 2016). And, in general, adults are better at solving analogies than children. But, when the required domain knowledge and a causal framing are present then children can solve analogies such as “body is to feet as table is to ?” as early as the 3-4 years-old (e.g., Goddu et al.,, 2020; Goswami,, 1991). And when analogies are presented in a more challenging or far context young children tend to revert to associative strategies, e.g., replying ’egg’ to ’dog is to doghouse as chicken is to ?’ instead of ’chicken coop’ (Stevenson and Hickendorff,, 2018; Gentner,, 1988; Thibaut and French,, 2016).
@@
There are many tasks used to study analogical reasoning and transfer in people, from verbal to geometric to scene analogy problems (e.g., Ichien et al.,, 2020; Richland et al.,, 2006; Mulholland et al.,, 1980). However, many of these tasks are either not suitable for children (e.g., verbal analogies may contain unfamiliar words or relations for children) or to LLMs (e.g., visual analogies designed for children are still difficult for today’s multimodal models (Yiu et al.,, 2024)). Therefore, we need a domain that is text-based, but doesn’t require domain knowledge beyond what a typical child or LLM would know. Letter-string analogies fit the bill as they require very little domain knowledge and offer an idealized scenario to examine analogical reasoning in a “pure, uncontaminated way” (Hofstadter,, 1984, p. 3). In these puzzles, a string of letters is transformed according to one or more rules, and the task is to use analogy and apply the same transformations to a new string. For example, “If abc changes to abd, what should pqr change to?” (Mitchell,, 2021).
@@
Letter-string analogy solving has been studied in human adults and LLMs. For example, Webb et al., (2023) showed that GPT-3 is able to solve letter-string analogies better than college students. Lewis and Mitchell, (2024) showed that GPT-models solved letter-string analogies at about 60% accuracy in the Latin alphabet domain, somewhat below the level of adults they tested. Interestingly, Lewis and Mitchell, (2024) and Hodel and West, (2023) found that GPT-3’s performance degraded when presented with these same analogies using an alphabet of shuffled letters. Moreover, Lewis and Mitchell, (2024) showed that GPT-models had great difficulty solving letter-string analogies in an unfamiliar alphabet of symbols, whereas people did not. As such, there is conflicting evidence of whether LLMs can generalize analogy solving to novel domains (Lewis and Mitchell,, 2024; Webb et al.,, 2024; Hodel and West,, 2023), something that comes easily to adults (e.g., Thibaut et al.,, 2022; Doumas et al.,, 2022), and that even children appear capable of when domains share structural similarities (Chen,, 1996; Gentner and Toupin,, 1986; Bobrowicz et al.,, 2020; Holyoak et al.,, 1984). Thus, while there is some evidence to suggest that LLMs can solve letter-string analogies at around the same level as people, it is unclear whether these models understand the problem and are actually using analogical reasoning (Opiełka et al.,, 2024; Stevenson et al.,, 2023; Moskvichev et al.,, 2023).
@@ 2 Method
We compared 42 children (7-9 year-olds), 62 adults, and 55 runs of each of four LLMs (Anthropic’s Claude-3.5, Google’s Gemma-2 27B, Open AI’s GPT-4o, and Meta’s Llama-3.1 405B) on a set of letter-string analogies under three alphabet conditions: Latin, Greek and a Symbol list.
@@ 2.1.1 Letter-String Analogy Task
Letter-string analogies, pioneered by Hofstadter 1984, are a type of analogy puzzle (A is to B as C is to D) involving alphabetic strings where one set of strings transforms to another, and the task is to use analogy to generalize the same transformation to a new string. For example, “If the string of letters abc changes to abd. How would you change the string pqrs in the ‘same way’?” (Mitchell,, 2021). Two things are happening here. First, the move from abc in term A to abd in term B shows that the last letter in the string, c, shifts to its successor in the alphabet, d. Second, the successor transformation must be generalized to the C term, a new string pqrt.
However, another possible (and more literal) solution to the problem could be pqrd. Here, we could apply a different rule, namely that the last letter is replaced with d. While letter-string analogies do not necessarily have a correct answer, there are answers that people tend to prefer, which is what we consider "correct" in this context. In this case, it would be pqrt.
All in all, there are several types of possible transformations from A to B and generalizations from A to C as described in (Webb et al.,, 2023). We use only the simplest transformations of successor, predecessor and repetition, and the generalizations are limited to shifting in the alphabet and letter repetitions, rules that children are expected to be familiar with.
The task items from the Latin alphabet are shown in Table 1.
Table 1: Letter-String Task, Latin Item Set
item A $\rightarrow$ B C $\rightarrow$ ? D AB Rule AC Rule
Latin 1 a b $\rightarrow$ a c g h $\rightarrow$ ? g i successor(last, 1) shift
Latin 2 c d $\rightarrow$ c c e e m n $\rightarrow$ ? m m o o successor(last, 1), repetition(all, 2) shift
Latin 3 e f $\rightarrow$ e h k l $\rightarrow$ ? k n successor(last, 2) shift
Latin 4 d e $\rightarrow$ d f f g h $\rightarrow$ ? g i i successor(last, 1), repetition(last, 2) shift
Latin 5 c d $\rightarrow$ b d m m n n $\rightarrow$ ? l l n n predecessor(first, 1), repetition(all, 2) shift, repetition(all, 2)
@@
Figure 1:
Visual representation of letter-string analogy item 1. (a) Latin alphabet serves as the baseline domain, (b) Greek alphabet as the near transfer domain, and (c) abstract symbols as the far transfer domain. Each panel shows the same analogical problem structure using different alphabets to test transfer.
Alphabets
For each of the items in the Latin alphabet we also created a near transfer version using the Greek alphabet and a far transfer version in our invented Symbol alphabet “*@ % ! ^ # ~ \$ { ? = :”, see Figure 1. We chose the Greek alphabet as near transfer domain because Greek symbols are somewhat visually similar to the Latin alphabet, but otherwise unfamiliar to the children in our study. We presented actual Greek symbols to humans, but chose the written version (i.e., alpha, beta, etc.) for LLMs based on their ability to list the Greek alphabet in this form upon request. We chose to use an ordered list of Symbols for far transfer, because it is an unfamiliar ‘alphabet’ that neither people nor the LLMs had seen before in this context, but at the same time were both able to process (i.e., the children can identify differences visually and for the LLMs these are common symbol keys on a keyboard). The constructed items for each alphabet were kept consistent, where the same transformations and generalizations from item 1 of the Latin alphabet were also used for item 1 of the Greek and Symbol alphabets, see Table 1 in the Appendix for an overview of all items.
@@ 2.1.2 Human Data Collection
Both children and adults completed the task in a browser. They were first shown the Latin alphabet and told that they would solve puzzles with these. For adults there was a simple example with feedback as the study was carried out fully online. For children, the interface was explained and demonstrated in person. Participants then solved two simple practice items without feedback (used to check understanding of task). Then for each alphabet, they were shown the list of letters/symbols and told they would again solve puzzles using these letters/symbols, where the Greek and Symbol alphabets were referred to as “secret code” letters for children. There were five items for each alphabet, with 15 items total. See Figure 2 for the instruction screens and Table 8 for an overview of all items.
(a) Initial instructions
(b) Latin alphabet instructions
(c) Greek/Symbol instructions
Figure 2:
Task instructions presented to adult participants. (a) General task introduction and practice instructions, (b) specific instructions for the Latin alphabet condition, and (c) modified instructions for Greek alphabet and Symbol conditions. These instruction screens were exclusive to adult participants and were not shown to children.
@@
We collected adult data online from fluent English speakers through Prolific online research participant recruitment platform. We recruited 68 adults of 18 years or older (M=24.0, SD =7.33, 50% female) who had an education level of completed secondary education or higher, and resided in the Netherlands or neighboring countries (as children were recruited in the Netherlands). We also required that they have no language disorders and have (corrected-to-) normal vision to ensure they could see/process the task, that they use a device at least 2x a week (to ensure digital fluency), and that they have a 95% or higher approval rating on Prolific to ensure high quality data from the participants. Based on the pre-registered exclusion criteria for adults (answering >80% of items), 6 adults were excluded.
@@
Children’s data was collected from 44 children (7-9 year-olds, M=8.26, SD=0.67) at a local school on an electronic tablet. The recruited school is a public Montessori school, located in an urban environment. The school emphasizes natural materials and does not use tablets or computers in the lower elementary groups. The researchers gave spoken instructions to the children given the limited reading abilities in this age group. The children were then allowed to complete the task independently. We excluded two of the 44 children, because they did not complete the task. We did not apply the pre-registered exclusion requirement of solving 40% of Latin alphabet items correctly, as the two children who made errors here accidentally submitted empty responses without making this mistake on other alphabets, and also met the other inclusion criteria.
@@ 2.1.3 LLM Data Collection
We collected data from LLMs from six types of models: Anthropic’s Claude-3.5 and Claude-3; Google’s Gemma-2-9B and Gemma-2 27B; Open AI’s GPT-3, GPT-3.5, GPT-4, and GPT-4o; Meta’s Llama-3.1-8B, Llama-3.1-70B, Llama-3.1-405B; Mistral AI’s Mixtral-8x7B and Mixtral-8x22B; Qwen’s Qwen-1.5-72B and Qwen-1.5-110B.
In order to keep the conditions for the LLM data collection as similar to that of people and, especially to fairly compare LLM results to those of children, we presented all analogies in a zero-shot setting using the same instructions that we spoke to the children, with the exception of referring to the Greek and Symbol alphabets as “secret code letters”. The LLMs received the same two practice items without feedback that the children received.
Please note that the alphabet was specified before each item, just like in the human data collection. Also, following Webb et al., (2023)’s approach to administering verbal analogies and digit matrices, all previous conversation with the LLM was pre-pended to each successive item so that the models could learn while testing just as people could. This seemed especially important because the exact same rules were applied in the same order from one alphabet set to the next. It is perhaps important to note that we also ran the tasks without pre-pending previous conversation, which generally resulted in lower LLM performance (see Appendix C).
We presented the analogies in chat completion mode using Python API’s from Anthropic for Claude models, from Open AI for GPT models and from Together AI for the remaining models, which are all open-source. We specified a temperature of 0 for near-deterministic data collection and set the maximum number of tokens to 10.
For each model type, the newest and largest model had the best performance. Furthermore, the Mistral and Qwen models performed far worse than the other models. Therefore, to provide clear and concise results we report the results based on this selected set of models: Anthropic’s Claude-3.5, Google’s Gemma-2 27B, Open AI’s GPT-4o, and Meta’s Llama-3.1 405B. The results of the other models can be found in the Supplementary Material on OSF.
@@
In order to be able to make robust comparisons between individual LLMs and groups of people, we adopted a similar methodology to Webb et al. 2023 and administered approximately as many variants of the task to each LLM as we had people who solved it. To do so, we created variants of each item by systematically shifting all of the characters in the item. For example, “a b : a c :: l m : ?” became “b c : b d :: m n : ?” and repeated this to create the required 54 total parallel testlets (based on power analysis for our pre-registration). This allowed us to have robust estimates of LLM performance, while creating some variation in the data and enabling us to compute SE’s for statistical analyses. Item difficulties for LLMs were expected to be very similar, but not exactly the same because LLM problem solving is sensitive to how often tokens occur in pre-training data (Razeghi et al.,, 2022), as can also be seen in the results by Webb et al. 2023.
Prompting templates
We administered each item using 5 different prompt templates, as prompt engineering can change the LLMs’ performance on the task. Examples of the templates and results can be found in Appendix 9. The template If a b c changes to a b d, what does i j k change to? worked best on the whole and was therefore chosen for LLM data collection.
Procedure
The LLM general instruction was as follows:
    We are going to do puzzles with the letters or symbols
    ‘a b c d e f g h i j k l m n o p q r s t u v w x y z’.
    Example
    if a changes to b, then j changes to k
This included a complete example that was not included in the human data collection.
The LLM item instruction was:
    The {letter|symbol} list is
    ‘{Latin alphabet|Greek alphabet|Symbol list}’.
    If {A} changes to {B}, what does {C} change to ?
The system instruction was:
    You are a helpful assistant that solves letter-string analogies.
    Only give the answer, no other words or text.

### v3

@@ Abstract
In people, the ability to solve analogies such as “body : feet :: table : ?” emerges in childhood, and appears to transfer easily to other domains, such as the visual domain “$($ : )) :: << : ?”. Recent research shows that large language models (LLMs) can solve various forms of analogies. However, can LLMs generalize analogy solving to other domains like people can? To investigate this, we had children, adults, and LLMs solve a series of letter-string analogies (e.g., a b : a c :: j k : ?) in the Latin alphabet, in a near transfer domain (Greek alphabet), and a far transfer domain (list of symbols). Children and adults easily generalized their knowledge to unfamiliar domains, whereas LLMs did not. This key difference between human and AI performance is evidence that these LLMs still struggle with robust human-like analogical transfer.
@@ 1 Introduction
You may be familiar with the analogy “consciousness is like an iceberg”. Here, people intuitively infer the below-the-surface depth and complexity of consciousness by relating it to an iceberg, whose mass is mostly found under water, just as our subconscious dwells under our conscious minds. This ability emerges in childhood (Goddu et al., 2020; Gentner, 1988; Stevenson and Hickendorff, 2018). However, it is a subject of debate whether analogical reasoning has emerged in Large Language Models (LLMs) (Webb et al., 2023; Lewis and Mitchell, 2025; Hodel and West, 2023; Webb et al., 2024). More importantly, are LLMs able to solve analogies at this level of conceptual abstraction and generalize to novel domains (Mitchell, 2021; Shiffrin and Mitchell, 2023)? In this study, we investigate analogical transfer at two levels of abstraction (near and far), and compare LLM performance not only to adults, but also to children, who are still developing analogical reasoning abilities. We ask the question: Can LLMs can generalize analogy solving like children can?
Figure 1: Human vs LLM performance on letter-string analogies across alphabet domains.
@@
Analogical reasoning, the process of applying a known concept to understand something new through relational similarity, is fundamental to the way people think and learn (Holyoak, 2012; Gentner and Hoyos, 2017). This is because we humans can easily generalize —that is, transfer principles discovered in one domain to new domains that share varying degrees of similarity with the original (Doumas et al., 2022). This can be principles in near contexts that are similar in terms of concrete attributes (e.g., shape, “a pyramid is like an iceberg”) or in farther contexts that are only similar in terms of abstract relations (e.g., abstraction of depth, “consciousness is like an iceberg”) (Barnett and Ceci, 2002). Near analogies tend to be easier for both adults and children to solve than far analogies (Johnson et al., 2025; Jones et al., 2022; Thibaut and French, 2016). And, in general, adults are better at solving analogies than children. But, when the required domain knowledge and a causal framing are present, then children can solve analogies such as “body is to feet as table is to ?” as early as the 3-4 years-old (e.g., Goddu et al., 2020; Goswami, 1991). And when analogies are presented in a more challenging or far context, young children tend to revert to associative strategies, e.g., replying ‘egg’ to ‘dog is to doghouse as chicken is to ?’ instead of ‘chicken coop’ (Stevenson and Hickendorff, 2018; Gentner, 1988; Thibaut and French, 2016).
@@
There are many tasks used to study analogical reasoning and transfer in people, from verbal to geometric to scene analogy problems (e.g., Ichien et al., 2020; Richland et al., 2006; Mulholland et al., 1980). However, many of these tasks are either not suitable for children (e.g., verbal analogies may contain unfamiliar words or relations for children) or to LLMs (e.g., visual analogies designed for children are still difficult for today’s multimodal models Yiu et al. (2024)). Therefore, we need a domain that is text-based, but doesn’t require domain knowledge beyond what a typical child or LLM would know. Letter-string analogies fit the bill as they require very little domain knowledge and offer an idealized scenario to examine analogical reasoning in a “pure, uncontaminated way” (Hofstadter, 1984, p. 3). In these puzzles, a string of letters is transformed according to one or more rules, and the task is to use analogy and apply the same transformations to a new string. For example, “If abc changes to abd, what should pqr change to?” (Mitchell, 2021).
@@
Letter-string analogy solving has been studied in human adults and LLMs. For example, Webb et al. (2023) showed that GPT-3 is able to solve letter-string analogies better than college students. Lewis and Mitchell (2025) showed that GPT-models solved letter-string analogies at about 60% accuracy in the Latin alphabet domain, somewhat below the level of adults they tested. Interestingly, Lewis and Mitchell (2025) and Hodel and West (2023) found that GPT-3’s performance degraded when presented with these same analogies using an alphabet of shuffled letters. Moreover, Lewis and Mitchell (2025) showed that GPT-models had great difficulty solving letter-string analogies in an unfamiliar alphabet of symbols, whereas people did not. As such, there is conflicting evidence of whether LLMs can generalize analogy solving to novel domains (Lewis and Mitchell, 2025; Webb et al., 2024; Hodel and West, 2023), something that comes easily to adults (e.g., Thibaut et al., 2022; Doumas et al., 2022), and that even children appear capable of when domains share structural similarities (Chen, 1996; Gentner and Toupin, 1986; Bobrowicz et al., 2020; Holyoak et al., 1984). Thus, while there is some evidence to suggest that LLMs can solve letter-string analogies at around the same level as people, it is unclear whether these models understand the problem and are actually using analogical reasoning (Opiełka et al., 2024; Johnson et al., 2025; Moskvichev et al., 2023).
@@ 2 Method
We compared 42 children (7-9 year-olds), 62 adults, and 55 runs of each of four LLMs (Claude-3.5 Anthropic (2024); Gemma-2 27B GemmaTeam (2024); GPT-4o OpenAI (2023); and Llama-3.1 405B Touvro et al (2023)) on a set of letter-string analogies under three alphabet conditions: Latin, Greek and a Symbol list.
@@ 2.1.1 Letter-String Analogy Task
Letter-string analogies, pioneered by Hofstadter 1984, are a type of proportional analogy (A is to B as C is to D) involving alphabetic strings. For example, “If the string of letters abc changes to abd. How would you change the string pqrs in the ‘same way’?” (Mitchell, 2021). Such letter-string analogies can be solved in multiple ways. For example, shifting the last letter and responding pqrt is what people tend to prefer (and what we consider “correct” in this context). However, another possible solution could be pqrd, where a literal rule is applied, namely replacing the last letter with d.
Rules
There are several possible transformations from A to B and generalizations from A to C as described in Webb et al. (2023). We use only the simplest transformations of successor (one and two after), predecessor (one before) and repetition, and the generalizations are limited to shifting in the alphabet and letter repetitions —rules that children are expected to be familiar with.
Alphabets
For each of the items in the Latin alphabet we also created a near transfer version using the Greek alphabet and a far transfer version in our invented Symbol alphabet: * @ % ! ^ # ~ \$ { ? = : (see Figure 2 for an example). We chose the Greek alphabet as near transfer domain because Greek symbols are somewhat visually similar to the Latin alphabet, but otherwise unfamiliar to the children in our study. We presented actual Greek symbols to humans, but chose the written version (i.e., alpha, beta, etc.) for LLMs based on their ability to list the Greek alphabet in this form upon request. We chose to use an ordered list of Symbols for far transfer, because it is an unfamiliar ‘alphabet’ that neither people nor the LLMs had seen before in this context, but at the same time were both able to process (i.e., the children can identify differences visually and for the LLMs these are common symbol keys on a keyboard). The constructed items for each alphabet were kept consistent, where the same transformations and generalizations from item 1 of the Latin alphabet were also used for item 1 of the Greek and Symbol alphabets. See Table 1 in the Results section for an overview of all items.
@@
Figure 2: Letter-string analogy task item 1 in (a) baseline alphabet: Latin, (b) near transfer alphabet: Greek, and (c) far transfer alphabet: Symbol.
@@ 2.1.2 Human Data Collection
Both children and adults completed the task in a browser. They were first shown the Latin alphabet and told that they would solve puzzles with these. For adults there was a simple example with feedback as the study was carried out fully online. For children, the interface was explained and demonstrated in person. Participants then solved two simple practice items without feedback (used to ensure understanding of task). Then for each alphabet, they were shown the list of letters/symbols and told they would again solve puzzles using these letters/symbols, where the Greek and Symbol alphabets were referred to as “secret code” letters for children. There were five items for each alphabet, with 15 items total.
@@
We collected adult data online from fluent English speakers through the Prolific research participant recruitment platform. We recruited 68 adults of 18 years or older (M=24.0, SD =7.33, 50% female) who had completed secondary education or higher and resided in the Netherlands or neighboring countries (as children were recruited in the Netherlands). We also required that they have no language disorders and have (corrected-to-) normal vision to ensure they could see/process the task, that they use a device at least 2x a week (to ensure digital fluency), and that they have a 95% or higher approval rating on Prolific to ensure high quality data from the participants. Based on the pre-registered exclusion criteria for adults (answering >80% of items), 6 adults were excluded.
@@
Data was collected from 44 children (7-9 year-olds, M=8.26, SD=0.67) at a local school on an electronic tablet. The recruited school is a public Montessori school and emphasizes natural materials and does not use tablets or computers in this age group. All children from the participating classrooms were included in the study, as language disorders are generally not yet assessed in this age group in the Netherlands. The researchers gave spoken instructions given the limited reading abilities in this age group. The children then completed the task independently. We excluded two children, because they did not complete the task.
@@ 2.1.3 LLM Data Collection
We collected data from LLMs from four types of models: Anthropic’s Claude-3 and Claude-3.5; Google’s Gemma-2-9B and Gemma-2 27B; Open AI’s GPT-3, GPT-3.5, GPT-4, and GPT-4o; Meta’s Llama-3.1-8B, Llama-3.1-70B, Llama-3.1-405B. For each model type, the newest and largest model had the best performance. Therefore, to provide clear and concise results our main results comparing human and LLM performance report on this selected set of models: Anthropic’s Claude-3.5, Google’s Gemma-2 27B, Open AI’s GPT-4o, and Meta’s Llama-3.1 405B. A brief overview of the results of other models can be found in Section 4; the full dataset is available in our GitHub repository.
Procedure
We presented the analogies in chat completion mode using Python APIs from Anthropic for Claude models, from Open AI for GPT models and from Together AI for the remaining models, which are all open source. We specified a temperature of 0 for near-deterministic data collection and set the maximum number of tokens to 10.
Prompt
The LLM general instruction was as follows: We are going to do puzzles with the letters or symbols ‘[Latin alphabet|Greek alphabet|Symbol list]’. Example ’if a changes to b, then j changes to k’.
Per item the LLMs received the instruction and item as follows: The [letter|symbol] list is ‘[Latin alphabet|Greek alphabet|Symbol list]’. If [A] changes to [B], what does [C] change to ?.
Pre-pending previous conversation
Also, following Webb et al. (2023)’s approach to administering verbal analogies and digit matrices, all previous conversation with the LLM was pre-pended to each successive item so that the models could learn while testing just as people could. This seemed especially important because the exact same items with the same rules were applied in the same order from one alphabet set to the next. We also ran the tasks without pre-pending previous conversation, which generally resulted in lower LLM performance (see Appendix C).
Prompting templates
We tested 5 different prompt templates for presenting items to LLMs, as prompt engineering can change the LLMs’ performance on the task. Results are reported for the best performing template as shown above: If A changes to B, what does C change to?. See the Appendix 8 for more details on the different templates and results.
Differences between Human and LLM Procedures
To keep the conditions for the LLM data collection similar to that of people and, especially to fairly compare LLM results to those of children, we presented all analogies in a zero-shot setting using the same instructions that we spoke to the children. There were two exceptions. First, with children we referred to the Greek and Symbol alphabets as “secret code letters”, whereas this was ‘alphabet’ and ‘(ordered) list’ respectively for adults and LLMs. Second, the LLMs received the worked example ’if a changes to b, then j changes to k’ that humans did not receive.
@@
To enable robust comparisons between individual LLMs and groups of people, we adopted a similar methodology to Webb et al. 2023 and administered approximately as many variants of the task to each LLM as we had people who solved it. To do so, we created variants of each item by systematically shifting all of the characters in the item. For example, “a b : a c :: l m : ?” became “b c : b d :: m n : ?”. For each of the 5 items per alphabet administered to humans (see Table 1), we created 4 item variants, i.e., shifted 1-4 elements to the left and/or right. We then systematically selected item variants to create 55 unique parallel testlets (required number based on power analysis from pre-registration; Note: we administered each testlet as a ‘conversation’ containing multiple messages to recreate how we tested humans, see 2.1.3). This allowed us to have robust estimates of LLM performance, while creating some variation in LLM data and enabling us to compute SE’s for statistical analyses.
Table 1: Base item set administered to adults, children, and LLMs.
Item ID Alphabet A B C D AB Rule AC Rule
A Practice a b j k successor_1 shift
B Practice c d c d d j k j k k repeat_1 shift
1 Latin a b a c g h g i successor_1 shift
2 Latin c d c c e e m n m m o o successor_1, repeat_2 shift
3 Latin e f e h k l k n successor_2 shift
4 Latin d e d f f g h g i i successor_1, repeat_1 shift
5 Latin c d b d m m n n l l n n predecessor_1 shift, repeat_2
1 Greek α β α γ ζ η ζ ϑ successor_1 shift
2 Greek γ δ γ γ ε ε ϰ λ ϰ ϰ µ µ successor_1, repeat_2 shift
3 Greek ε ζ ε ϑ ι ϰ ι µ successor_2 shift
4 Greek η ϑ η ι ι λ µ λ ν ν successor_1, repeat_1 shift
5 Greek β γ α γ ν ν ξ ξ µ µ ξ ξ predecessor_1 shift, repeat_2
1 Symbol * @ * % ~ \$ ~ { successor_1 shift
2 Symbol % ! % % ˆ ˆ = : = = ) ) successor_1, repeat_2 shift
3 Symbol @ % @ ˆ # ~ # { successor_1 shift
4 Symbol ! ˆ ! # # \$ { \$ = = successor_1, repeat_1 shift
5 Symbol ˆ # ! # = = : : { { : : predecessor_1 shift, repeat_2